OpenStudy (anonymous):

A fruit company delivers its fruit in two types of boxes: large and small. A delivery of 7 large boxes and 9 small boxes has a total weight of 192 kilograms. A delivery of 5 large boxes and 3 small boxes has a total weight of 114 kilograms. How much does each type of box weigh?

OpenStudy (anonymous):

then a would =75/4 b=27/4